Can you answer this question (and try to show your steps, if possible). Thank you.
Given f(x) = 2x^2+8x-4 and g(x)=-5x+6, find (f-g)(x)

Answer by Nfrey78(18)

Can you answer this question (and try to show your steps, if possible). Thank you.
Given f(x) = 2x^2+8x-4 and g(x)=-5x+6, find (f-g)(x)
People get most confused with the notation.
(f-g)(x) = f(x)-g(x)
This notation just means we are subtracting the two functions to create a new function.
(2x^2 + 8x - 4) - (-5x + 6)
Since we are subtracting that second polynomial make sure to distribute the negative, i.e. flipping the signs in the second expression to get
2x^2 + 8x - 4 + 5x - 6
Simplify by combining like terms
2x^2 + 13x -10
So (f - g)(x) = 2x^2 + 13x -10